OpenStudy (thebossofme191):

A survey shows 25 people like pepperoni pizza and 31 like banana pepper pizza and 5 like both if 66 people were surveyed how many didn't like banana pepper pizza a.20 b.34 c.31 d.35

OpenStudy (thebossofme191):

@escamer

OpenStudy (thebossofme191):

@pooja195

OpenStudy (purple_pink):

do the math

OpenStudy (purple_pink):

66-31=?

OpenStudy (purple_pink):

wait hold on no

OpenStudy (bemoseemo):

66-25=?

OpenStudy (purple_pink):

http://prntscr.com/a1luy9

OpenStudy (purple_pink):

31+5=?

OpenStudy (purple_pink):

@Thebossofme191 are you there?

OpenStudy (thebossofme191):

@purple_pink yeah thx